Batman104:

One brand of vinegar has a pH of 4.5. Another brand has a pH of 5.0. The equation for the pH of a substance is pH = –log[H+], where H+ is the concentration of hydrogen ions. What is the approximate difference in the concentration of hydrogen ions between the two brands of vinegar? 2.2 *10^-6 3.2*10^-1 3.2*10^1 6.8*10^4

supie:

I don't really know this stuff but I think you have to do \[5 = -log H+\]\[5 = log (1 / H+)\]\[1/ H+ = 10^5\]\[H+ = 10^-5\]

Batman104:

thanks

supie:

np